# Finite Distance Corrections to Vacuum Birefringence in Strong Gravitational and Electromagnetic Fields

### v1

@@ V.2.3 Comparison of QED and Born-Infeld signatures
We have seen that the perturbative formalism successfully isolates the distance-dependent reduction of birefringence. For both QED and Born-Infeld theories in a monopole background, the signal is suppressed by a $(b/r)^{4}$ factor for nearby sources. This steep dependence implies that X-ray polarimetry of sources at finite distances (such as binary partners or surface emission) must account for these truncation factors to avoid underestimating the intrinsic non-linear coupling parameters.
@@ V.3 The magnetar observer test
The theoretical distinctions between QED and Born-Infeld birefringence derived in the preceding subsections find their most critical application in the study of magnetars, which are neutron stars possessing ultra-strong magnetic fields ($B\sim 10^{14}-10^{15}$ G). In these environments, the effective NLED coupling is enhanced by the extreme field strength, making vacuum birefringence potentially observable. However, current models often rely on asymptotic approximations that assume the photon traverses the entire magnetic dipole field from infinity to infinity (or from the center to infinity). Here, we demonstrate that the finite-distance corrections derived in Eq. (31) and Eq. (35) introduce a significant suppression factor for surface emission, particularly for radiation emerging from the stellar limb.
@@ V.3.1 Geometric configuration and suppression factor
We consider a physical scenario where polarized X-rays are emitted from a hot spot on the surface of a magnetar with radius $R_{NS}$ and observed by a distant telescope ($r_{\text{R}}\to\infty$). In this limit, the term depending on the observer’s position in Eq. (35) vanishes ($c_{\text{R}}(b/r_{\text{R}})^{4}\to 0$). The source term, however, remains dominant.
@@
For a physical magnetar characterized by a dipole field, the background magnetic field decays as $B\propto r^{-3}$, leading to an energy density scaling of $r^{-6}$. Consequently, the finite-distance suppression factor is significantly steeper than the monopole approximation ($n=4$) derived in the previous subsection. Adapting the general series expansion (Eq. (31)) to the dipole case ($n=6$), the observable differential deflection angle becomes:
$\Delta\alpha_{obs}=\Delta\alpha_{\infty}\left[1-c_{\text{S}}\left(\frac{R_{NS}\sin\psi}{R_{NS}}\right)^{6}\right].$ (42)
For surface emission, the photon originates at the periastron of the would-be scattering orbit and travels only outward. This corresponds to exactly half of the scattering path for a symmetric field distribution. Utilizing this symmetry, the coefficient for surface emission is exactly $c_{\text{S}}=1/2$ relative to the full scattering path. The suppression profile is:
$\boxed{\mathcal{F}_{\text{dipole}}(\psi)=1-\frac{1}{2}\sin^{6}\psi},$ (43)
which implies that for limb emission ($\psi\to\pi/2$), the observed birefringence is exactly half the value predicted by asymptotic scattering models.
@@ V.3.2 Observational consequences for X-ray polarimetry
The implications of Eq. (43) are substantial for the interpretation of data from X-ray polarimetry missions. Standard analyses often assume that the birefringence accumulates over a characteristic length scale $L\sim R_{NS}$, yielding a total phase shift proportional to the full path integral. Our derivation shows that this assumption leads to a systematic overestimation of the NLED coupling strength if the finite-distance truncation is ignored.
Specifically, for a magnetic dipole background where $n=6$ (as discussed in Eq. (40) for the dipole case), the suppression is even more severe at large angles:
$\mathcal{F}_{\text{dipole}}(\psi)\approx 1-\tilde{c}_{\text{S}}\sin^{6}\psi.$ (44)
Because the field decays rapidly ($B\propto r^{-3}$), the bulk of the birefringence in the asymptotic model is generated in the immediate vicinity of the perigee $r\approx b$. For surface emission, the photon originates at this region of maximum interaction but travels only outward. Consequently, it effectively experiences only half of the lensing potential (or less, depending on $\psi$) compared to a photon passing the star from behind.
To visualize the magnitude of this effect, we plot the theoretical suppression profile derived in Eq. (43) in Figure 2. We compare the physical dipole scaling ($n=6$) relevant for magnetars against the monopole approximation ($n=4$).
Figure 2: The Finite-Distance Suppression Factor $\mathcal{F}(\psi)$ as a function of the emission angle $\psi$. The solid black line represents the physical model for a magnetar (dipole scaling, $n=6$), while the dashed red line shows the monopole approximation ($n=4$). For grazing emission ($\psi\to 90^{\circ}$), the observable birefringence is suppressed by exactly 50% ($c_{\text{S}}=1/2$) due to the truncation of the photon path relative to the asymptotic scattering limit.
Figure 3 explicitly tracks the accumulation of the birefringence along the line of sight. For the dipole case (solid black line), we observe a steep step-function behavior: approximately $80\%$ of the total signal is generated within a narrow angular window of $\pm 30^{\circ}$ around the periastron (the point of closest approach, $\phi=90^{\circ}$). This confirms that the NLED interaction is confined to a compact zone of influence near the stellar surface. Consequently, when a photon is emitted from the surface (truncating the path at $90^{\circ}$), it fundamentally misses half of this interaction zone, resulting in the drastic $50\%$ suppression factor derived in Eq. (43). This localization justifies the use of the steep power-law correction series in Eq. (31), and underscores the necessity of the geometric calibration for surface polarimetry.
Figure 3: The cumulative fraction of the total birefringence signal accumulated along the photon path angle $\phi$, where $0^{\circ}$ corresponds to the observer and $180^{\circ}$ to a source at asymptotic infinity. The periastron (point of closest approach to the lens) is located at $90^{\circ}$. For the magnetar dipole field (solid black line), the signal accumulation is highly localized, with the majority of the phase shift generated in the immediate vicinity of the periastron. This localization explains why surface emission (which effectively starts the integration at $90^{\circ}$) suffers a substantial geometric suppression relative to the full scattering path.
@@ V.3.3 Constraints on physical parameters
The suppression plotted above indicates that at high emission angles (limb emission), the observable birefringence may be reduced by up to 50% relative to the standard full-path integration (assuming $c_{\text{S}}\approx 0.5$ due to the half-path geometry).
Such a correction must be incorporated into the vacuum resonance models used to constrain the neutron star surface magnetic field $B_{\text{surf}}$ and the QED coupling $\beta_{\text{QED}}$. If the finite-distance correction is neglected, an observer measuring a specific polarization angle would incorrectly infer a weaker magnetic field or a smaller NLED coupling constant than actually exists. Therefore, precise tests of QED in the strong-field regime requires the use of the corrected estimator:
$\Delta\alpha_{\text{true}}\approx\frac{\Delta\alpha_{\text{measured}}}{\mathcal{F}(\psi)}.$ (45)
By applying this correction, future analyses of magnetar polarization data can decouple the geometric finite-distance effects from genuine deviations in the fundamental QED Lagrangian (such as those predicted by Born-Infeld theory in Eq. (39)). It ensures that any observed anomaly is attributed correctly to the NLED physics rather than the proximity of the source.
@@ VI Conclusion
In this work, we have developed a rigorous perturbative framework to calculate the finite-distance corrections to vacuum birefringence in strong gravitational and electromagnetic fields. By extending the geometric optical formalism to include the specific case of sources and observers located at finite radial coordinates, we addressed a critical gap in the theoretical modeling of non-linear electrodynamics (NLED) effects near compact objects.
Our analysis began with the linearization of the effective optical metric for weak NLED couplings, isolating the specific contributions of the photon’s polarization to the geodesic structure. A key result of our study is the derivation of the generalized differential deflection angle, Eq. (28), which explicitly separates the refractive curvature effects from the path-induced geometric shifts. We demonstrated that even in spherically symmetric backgrounds, the splitting of photon trajectories for different polarization modes generates a non-negligible geodesic shift contribution to the observable birefringence.
Applying this formalism to physical scenarios, we computed the explicit finite-distance correction series for both Quantum Electrodynamics (Euler-Heisenberg) and Born-Infeld theory. We found that for a magnetic monopole background, the finite-distance correction scales as $(b/r)^{4}$. The steep radial dependence signifies that vacuum birefringence is a highly localized phenomenon, with the bulk of the phase shift accumulating in the immediate vicinity of the source.
Crucially, our application to magnetar physics in Section V.3 revealed that standard asymptotic approximations significantly overestimate the observable polarization rotation for surface emission. We derived a geometric suppression factor, $\mathcal{F}_{\text{dipole}}(\psi)\approx 1-\frac{1}{2}\sin^{6}\psi$, which shows that for radiation emitted from the stellar limb, the observable effect is reduced by approximately 50% relative to full-scattering models. The correction is essential for the accurate interpretation of data from current and future X-ray polarimetry missions, ensuring that constraints on the surface magnetic field and the QED coupling constant are not systematically biased by geometric truncation.

### v2

@@ V.2.3 Multipole sensitivity of finite-distance suppression
We have seen that the perturbative formalism successfully isolates the distance-dependent reduction of birefringence. For the monopole benchmark (with $n=4$) the signal is suppressed by a $(b/r)^{4}$ factor for nearby sources, while for dipole-like scaling ($n=6$) the convergence to the asymptotic limit is even faster. This steep dependence implies that X-ray polarimetry of sources at finite distances (such as binary partners or surface emission) must account for these truncation factors to avoid underestimating the intrinsic non-linear coupling parameters.
@@ V.3 The magnetar observer test
The finite-distance corrections derived above find their most critical application in the study of magnetars, neutron stars with ultra-strong magnetic fields ($\mathcal{B}\sim 10^{14}--10^{15}\,\mathrm{G}$). We emphasize, however, that the Gauss-Bonnet construction in Secs. III-IV assumes an effective optical metric of static, spherically symmetric (SSS) form. A realistic magnetar magnetosphere is dipole-dominated and therefore axisymmetric [48, 50], so an exact dipole calculation lies outside the present SSS framework. Accordingly, in this subsection we use the dipole only to extract the radial falloff exponent controlling the truncation series and to illustrate the purely geometric suppression produced by finite-distance emission, particularly for near-limb rays.
@@ V.3.1 Geometric configuration and suppression factor
We consider a physical scenario where polarized X-rays are emitted from a hot spot on the surface of a magnetar with radius $R_{NS}$ and observed by a distant telescope ($r_{\text{R}}\to\infty$). In this limit, the term depending on the observer’s position in Eq. (IV.12) vanishes ($c_{R}(b/r_{R})^{n}\to 0$). The source term, however, remains dominant.
@@
For a physical magnetar characterized by a dipole field, the background magnetic field decays as $\mathcal{B}\propto r^{-3}$, leading to an energy density scaling of $r^{-6}$. Consequently, the finite-distance suppression factor is significantly steeper than the monopole approximation ($n=4$) derived in the previous subsection. Adapting the general series expansion (Eq. (IV.12)) to the dipole case ($n=6$), the observable differential deflection angle becomes:
$\Delta\alpha_{obs}=\Delta\alpha_{\infty}\left[1-c_{\text{S}}\left(\frac{R_{NS}\sin\psi}{R_{NS}}\right)^{6}\right].$ (V.18)
For surface emission, the photon originates near the periastron of the would-be scattering orbit and travels only outward. If one compares this one-sided trajectory to a two-sided scattering orbit through an idealized symmetric extension of the exterior field, the missing inbound segment suggests a truncation coefficient of order unity; in the perfectly symmetric benchmark one obtains $c_{S}=1/2$. We therefore adopt $c_{S}\simeq 1/2$ as an illustrative reference value, while stressing that for a realistic axisymmetric dipole magnetosphere (and for non-equatorial emission) the effective $c_{S}$ can differ by an $\mathcal{O}(1)$ factor.
With $b\simeq R_{\rm NS}\sin\psi$ this yields the benchmark suppression profile
$F_{\rm dipole}(\psi)\equiv\frac{\Delta\alpha_{\rm obs}}{\Delta\alpha_{\infty}}\simeq 1-\frac{1}{2}\sin^{6}\psi,$ (V.19)
which indicates an order-unity reduction for near-limb emission within this symmetric one-sided benchmark.
Figure 1: Finite-distance suppression factor for surface emission under benchmark comparisons. We plot the geometric reduction $F(\psi)\equiv\Delta\alpha_{\rm obs}/\Delta\alpha_{\infty}$ versus emission angle $\psi$. The dashed line shows the asymptotic two-sided scattering normalization ($F=1$). The solid black curve is the dipole-like benchmark $F_{\rm dipole}(\psi)\simeq 1-\tfrac{1}{2}\sin^{6}\psi$ (Eq. (V.19)). The dotted black curve shows a more aggressive one-sided benchmark with $c_{S}=1$, while the blue curve illustrates the monopole-like exponent ($n=4$) at the same reference coefficient $c_{S}=\tfrac{1}{2}$, highlighting the sensitivity of the angular dependence to the far-field decay exponent $n$.
Figure 1 makes explicit what “suppression” means by comparing the one-sided surface-emission geometry to the two-sided asymptotic scattering normalization ($F=1$). The dipole-like benchmark (solid black) shows that the truncation is negligible for near-radial emission but becomes rapidly important toward the limb due to the steep $\sin^{6}\psi$ dependence. The dotted curve ($c_{S}=1$) provides an upper benchmark for how strong the suppression could be under a more extreme one-sided mapping, while the monopole-like comparison (blue, $n=4$) demonstrates that the exponent controlling the finite-distance series directly governs how quickly the suppression turns on with emission angle. This overlay therefore separates the robust geometric trend (strongest effect near the limb) from the model-dependent normalization encoded in $c_{S}$ and from the field-structure dependence encoded in $n$.
@@ V.3.2 Observational consequences for X-ray polarimetry
The implications of Eq. (V.19) are substantial for the interpretation of data from X-ray polarimetry missions. Standard analyses often assume that the birefringence accumulates over a characteristic length scale $L\sim R_{NS}$, yielding a total phase shift proportional to the full path integral. Our derivation shows that this assumption leads to a systematic overestimation of the NLED coupling strength if the finite-distance truncation is ignored.
More generally, any dipole-like falloff leading to $n=6$ (see Eq. (V.16)) suggests a one-sided suppression of the form
$F_{\rm dipole}(\psi)\simeq 1-c_{S}\sin^{6}\psi,$ (V.20)
with $c_{S}=\mathcal{O}(1)$ set by the detailed (generally axisymmetric) geometry and by the mapping between the physical emission problem and the two-sided scattering normalization used to define $\Delta\alpha_{\infty}$. Because the field decays rapidly ($B\propto r^{-3}$), the bulk of the birefringence in the asymptotic model is generated in the immediate vicinity of the perigee $r\approx b$. For surface emission, the photon originates at this region of maximum interaction but travels only outward. Consequently, it effectively experiences only half of the lensing potential (or less, depending on $\psi$) compared to a photon passing the star from behind.
@@ V.3.3 Constraints on physical parameters
The benchmark suppression above indicates that at high emission angles (near-limb emission) the observable birefringence can be reduced at the order-tens-of-percent level relative to a two-sided full-path integration; in the symmetric one-sided benchmark ($c_{S}=1/2$) the reduction approaches $\sim 50\%$ as $\psi\to\pi/2$. Such geometric truncation must be accounted for when using surface-emission polarimetry to constrain the neutron star magnetic field and the effective QED/NLED coupling, otherwise the inferred parameters will be systematically biased. Accordingly, any inversion of data should use a model-dependent geometric factor $F(\psi)$:
$\Delta\alpha_{\rm true}\approx\frac{\Delta\alpha_{\rm measured}}{F(\psi)}\quad\text{(model-dependent)}.$ (V.21)
By applying this correction, future analyses of magnetar polarization data can decouple the geometric finite-distance effects from genuine deviations in the underlying birefringent vacuum response (e.g. higher-order QED contributions or other birefringent NLED corrections). It ensures that any observed anomaly is attributed correctly to the NLED physics rather than the proximity of the source.
@@ VI Conclusion
In this work, we have developed a rigorous perturbative framework to calculate the finite-distance corrections to vacuum birefringence in strong gravitational and electromagnetic fields. By extending the geometric optical formalism to include the specific case of sources and observers located at finite radial coordinates, we addressed a critical gap in the theoretical modeling of NLED effects near compact objects.
Our analysis began with the linearization of the effective optical metric for weak NLED couplings, isolating the specific contributions of the photon’s polarization to the geodesic structure. A key result of our study is the derivation of the generalized differential deflection angle, Eq. (IV.8), which explicitly separates the refractive curvature effects from the path-induced geometric shifts. We demonstrated that even in spherically symmetric backgrounds, the splitting of photon trajectories for different polarization modes generates a non-negligible geodesic shift contribution to the observable birefringence.
Applying this formalism to physical scenarios, we computed the explicit finite-distance correction series for Quantum Electrodynamics (Euler-Heisenberg) in a magnetic monopole benchmark and found that the leading finite-distance correction scales as $(b/r)^{4}$. The steep radial dependence signifies that vacuum birefringence is a highly localized phenomenon, with the bulk of the phase shift accumulating in the immediate vicinity of the source. As a consistency check, we also revisited Born-Infeld electrodynamics and confirmed that it is non-birefringent in the present sense: $\Delta\alpha\equiv\alpha_{+}-\alpha_{-}$ vanishes identically [13, 67].
Crucially, our application to magnetar physics in Section V.3 revealed that standard asymptotic approximations significantly overestimate the observable polarization rotation for surface emission. Crucially, our magnetar-motivated application in Sec. V.3 shows that standard asymptotic (two-sided) approximations can overestimate the observable polarization-dependent deflection for surface emission because the physical trajectory is one-sided. Within a symmetric benchmark mapping this yields a suppression profile $F_{\rm dipole}(\psi)\simeq 1-\tfrac{1}{2}\sin^{6}\psi$, indicating order-unity reductions for near-limb rays. Since a realistic dipole magnetosphere is axisymmetric rather than spherically symmetric, this should be interpreted as an illustrative scaling estimate; a fully axisymmetric extension is left for future work.